OpenStudy (anonymous):

a bag contains only red, yellow and blue discs. the probability of drawing a red disc at random is 1/2 the probability of drawing a yellow disc at random is x the probability of drawing a blue disc at random is 4x one disc is to be selected at random. work out the probability that it will be a blue disc. give your answer as a numerical value. could someone help me please:)

OpenStudy (anonymous):

is it 4/10 or 0.40 or 40% ?

OpenStudy (cwrw238):

one or the other

OpenStudy (mertsj):

\[\frac{1}{2}+x+4x=1\] \[5x=\frac{1}{2}\] \[x=\frac{1}{10}\]

OpenStudy (mertsj):

Probability of blue = 4/10=2/5